OpenStudy (anonymous):

How many grams of gold (Au) are there in 100 million atoms of Au?

OpenStudy (unklerhaukus):

\[M_{\text{Au}}=196[\text{g}/\text{mole}]\] \[1[\text{mole}]=6.022\times10^{23} [\text{atoms}]\] \[100 \times10^6 [\text{atoms}]\]\[=100 \times10^6 [\text{atoms}]\times\frac{1[\text{mole}]}{6.022\times10^{23} [\text{atoms}]}\times196[\text{g}/\text{mole}]\]\[=\quad\dots[\text g]\]